# AIR-Bench: Automated Heterogeneous Information Retrieval Benchmark

### v3

@@ Abstract
Jianlyu Chen1,2,6   Nan Wang3   Chaofan Li2,4   Bo Wang3 Shitao Xiao2   Han Xiao3   Hao Liao5   Defu Lian1,6   Zheng Liu2thanks: Corresponding authors 1University of Science and Technology of China 2Beijing Academy of Artificial Intelligence    3Jina AI 4Beijing University of Posts and Telecommunications    5Shenzhen University 6State Key Laboratory of Cognitive Intelligence chenjianlv@mail.ustc.edu.cn    research@jina.ai liandefu@ustc.edu.cn    zhengliu1026@gmail.com
@@ 2.4 Quality Control
3 Overview of AIR-Bench
In this section, we provide an overview of AIR-Bench from three perspectives: tasks, datasets and software.
Datasets. As shown in Figure 2, AIR-Bench currently has two distinct versions, 24.04 and 24.05, where the latest version 24.05 consists of a total of 69 datasets, covering 9 domains and 13 languages on two retrieval tasks. We hope to incorporate more domains and languages in the future version to provide an increasingly comprehensive evaluation benchmark for community developers. More details are available in Appendix B.1.
4 Diversity Analysis
@@
To analyze the query type diversity of AIR-Bench, we utilize GPT-4o444 gpt-4o-2024-08-06 Achiam et al. (2023) as labeler to label the type of the generated queries. Specifically, given a query, we prompt GPT-4o to select the most suitable type for the query from the optional types.
The statistics are grouped by tasks and splits in Table 3. Based on the results, we can make the following observations. Firstly, both the QA and Long-Doc tasks have the highest frequency of what queries, followed by claim queries as the second most common, and how queries as the third. Additionally, the QA task exhibits a more balanced distribution of the other query types, whereas the Long-Doc task shows a lower frequency of when queries and where queries. Lastly, a small number of queries are classified as others, reflecting the diverse types of queries present in AIR-Bench to some extent. Further diversity analysis of AIR-Bench is presented in Appendix B.3.
@@ 5 Experiment
RQ2: What additional evaluation functionalities dose AIR-Bench offer compared to MTEB/BEIR?
@@ 5.1 Consistency Analysis (RQ1)
#corpus #queries #positives
R-MSMARCO 8,841,823 6,980 7,437
G-MSMARCO 8,872,840 6,319 31,447
      w/o quality control 8,878,865 7,429 7,429
@@ 5.2 Comparison with MTEB/BEIR (RQ2)
Main Results. As presented in Table 5, we can make the following observations based on the comparison results. 1) LLM-based embedding models generally outperform large-size embedding models on both AIR-Bench and MTEB/BEIR, largely due to the superior generalization ability of LLMs. Besides, BM25 performs worse than all embedding models on both AIR-Bench and BEIR. 2) The QA task and the Long-Doc task in AIR-Bench exhibit a level of heterogeneity. The Spearman rank correlation coefficient between the rankings of the nine LLM-based embedding models across the two tasks is only 0.6, with a p-value of 0.0876. Moreover, as a large-size embedding model, e5-large-v2 even outperforms some LLM-based embedding models on the Long-Doc task. 3) By comparing the results on AIR-Bench and MTEB/BEIR, we observe that better performance on MTEB/BEIR may not indicate better performance on AIR-Bench. For example, according to Li et al. (2024), bge-en-icl utilizes more in-domain training data in MTEB/BEIR than bge-en-icl-e5data and achieves more superior performance on MTEB/BEIR. However, compared to bge-en-icl-e5data, bge-en-icl shows performance degradation on AIR-Bench, including both the QA task (54.46 $\rightarrow$ 53.60) and the Long-Doc task (73.43 $\rightarrow$ 72.62). This suggests that increased in-domain training data in MTEB/BEIR may introduce a risk of over-fitting, leading to a decline in the generalization ability of embedding models.
@@ 6.1 Evaluation Datasets for IR
In recent years, a series of milestone works have been introduced to the community. As the earlier contributions, MS MARCO Bajaj et al. (2016) includes Bing search questions paired with human-labeled relevant passages from Web documents. Natural Questions (NQ) Kwiatkowski et al. (2019) consists of Google search queries with human-labeled relevant Wikipedia pages. Both MS MARCO and NQ are designed for open-domain question answering tasks in English. Recent works like Mr.TyDi Zhang et al. (2021), while MIRACL Zhang et al. (2023) focus on multilingual retrieval in non-English languages. Mr.TyDi covers 11 languages and MIRACL encompasses an extended 18 languages. BEIR Thakur et al. (2021) and MTEB Muennighoff et al. (2023) are introduced to benchmark IR models in a general-domain zero-shot setting, including multiple existing datasets from diverse tasks and domains.
@@ 7 Conclusion
While AIR-Bench aims to be a comprehensive IR benchmark, incorporating new features to address limitations in existing benchmarks, it still has several inherent constraints: 1) Dependence on real-world corpora. The generation process for each dataset in AIR-Bench begins with corpora preparation. Consequently, the availability of real-world corpora is essential for generating an evaluation dataset. Future extensions of AIR-Bench will be significantly influenced by this factor. 2) Reliance on capabilities of LLM. We design complex prompts for the LLM to diversify the generated queries. The quality of these queries is largely dependent on the LLM’s capabilities, including its instruction-following, domain-specific, and language-specific capabilities. 3) Potential biases from quality control models. In addition to the LLM, we employ several existing IR models during the quality control stage. This reliance may introduce potential biases in the final evaluation dataset.
@@ Appendix A Details on Benchmark Construction
In this section, we provide more details on the construction of datasets in AIR-Bench. Considering the powerful capability of GPT-4 Achiam et al. (2023), we use gpt-4-1106-preview [^10] https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 as the LLM through the generation pipeline.
@@ A.1 Corpora Preparation
AIR-Bench currently covers two different tasks: QA and Long-Doc. For QA task, we directly use the real-world dataset as the corpus, such as Wikipedia, mC4Raffel et al. (2020), CC-NewsHamborg et al. (2017), etc. We filter out text that is either too short or too long and make a straightforward attempt to remove any information that names or uniquely identifies individuals, as well as any offensive content. For Long-Doc task, we first select one long document for each dataset, such as book, ArXiv paper, legal document, etc., and remove table of contents and references. Then we use the node parser [^11] SimpleNodeParser: https://github.com/run-llama/llama_index tool from LlamaIndexLiu (2022) to split the long document into fixed-size chunks121212chunk_size=200, chunk_overlap=50 as the corpus. All corpora used in AIR-Bench are available in Appendix B.1.
@@ A.3 Quality Control
Embedding Model. Considering that the corpora in AIR-Bench are multilingual, we use bge-m3131313https://huggingface.co/BAAI/bge-m3 as the embedding model to recall the top-1000 relevant documents.
Multiple Re-ranking Models. For the datasets in English and Chinese, we use the following three re-ranking models: bge-reranker-large [^14] https://huggingface.co/BAAI/bge-reranker-large, bce-reranker-base_v1151515https://huggingface.co/maidalun1020/bce-reranker-base_v1, mmarco-mMiniLMv2-L12-H384-v1161616https://huggingface.co/nreimers/mmarco-mMiniLMv2-L12-H384-v1. For the datasets in the other languages, we use the following three re-ranking models: bge-reranker-v2-m3171717https://huggingface.co/BAAI/bge-reranker-v2-m3, mmarco-mMiniLMv2-L12-H384-v1, bge-reranker-v2-gemma [^18] https://huggingface.co/BAAI/bge-reranker-v2-gemma.
@@ 24.05 Oct 17, 2024 9 13 69 Table 17-21
For each dataset, we use the same format as BEIR, i.e. corpus, queries and qrels191919qrels are the relevance labels for queries. The relevance label is 1 for the positive document, and 0 for the negative document., which are all available in the Hugging Face Hub [^20] https://huggingface.co/AIR-Bench of AIR-Bench. To avoid the possible data leakage, we keep the qrels in test splits private. For the qrels in dev splits, we make them public to enable the developers to perform evaluation by themselves.
@@ B.2 Licenses
In Table 16-21, we also list the licenses of the source corpora used for the dataset generation in AIR-Bench. All generated testing data in AIR-Bench is licensed under CC BY-NC-SA-4.0232323https://creativecommons.org/licenses/by-nc-sa/4.0. The testing data in AIR-Bench may only be used for evaluation purposes.
@@ B.3.1 Query Diversity
We also analyze the style diversity of the generated queries in AIR-Bench. We still utilize GPT-4o242424 gpt-4o-2024-08-06: https://platform.openai.com/docs/models/gpt-4o as labeler to label the style of the queries in AIR-Bench. The optional query styles include: formal, informal, professional, casual, complicated, concise, academic, and others. The statistics are grouped by tasks and splits in Table 10.
@@ Appendix C AIR-Bench Software
Appendix D Examples

### v4

@@ Abstract
Jianlyu Chen1,2,6   Nan Wang3   Chaofan Li2,4   Bo Wang3 Shitao Xiao2   Han Xiao3   Hao Liao5   Defu Lian1,6   Zheng Liu2,7thanks: Corresponding authors 1University of Science and Technology of China 2Beijing Academy of Artificial Intelligence    3Jina AI 4Beijing University of Posts and Telecommunications    5Shenzhen University 6State Key Laboratory of Cognitive Intelligence   7Hong Kong Polytechnic University chenjianlv@mail.ustc.edu.cn   research@jina.ai,    haoliao@szu.edu.cn liandefu@ustc.edu.cn    zhengliu1026@gmail.com
@@ 2.4 Quality Control
2.5 Design Motivations
We elaborate the design motivations of the data generation pipeline of AIR-Bench as follows.
Reliance on real-world corpora. Real-world corpora are usually diverse and available. Generating testing data based on real-world corpora not only closely aligns with real-world scenarios, but also significantly reduces the generation cost.
Generation of characters and scenarios. First, this step brings more transparency and interpretability on how a query is generated, compared to the naive method which directly prompts LLMs for query generation. Second, the generation of character and scenario also leads to higher diversity of queries, which contributes to the comprehensiveness of evaluation.
Query Rewriting. Through rewriting, queries are transformed into different forms while retaining equivalent semantics, which significantly increases the difficulty of retrieval tasks.
Generation of hard negatives. Similar to the introduction of query rewriting, this step increases the hardness of evaluation.
Quality Control. This step helps to remove low-quality queries and correct false relevance labels. Similar operations were also conducted in previous benchmark, e.g., the relevance assessment phase in MIRACL Zhang et al. (2023).
3 The AIR-Bench Benchmark
3.1 Overview
@@
LLM for Generation. We use powerful GPT-4222gpt-4-1106-preview: https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 Achiam et al. (2023) as the LLM through the generation pipeline. When prompting GPT-4, we set the sampling temperature to 1.0 to encourage more diversity.
Datasets. As shown in Figure 2, AIR-Bench currently has two distinct versions, 24.04 and 24.05, where the latest version 24.05 consists of a total of 69 datasets, covering 9 domains3339 domains: News, Web, Wiki, Science, Finance, Healthcare, Law, ArXiv, Book. and 13 languages44413 languages: English, Chinese, Spanish, French, German, Russian, Japanese, Korean, Arabic, Persian, Indonesian, Hindi, Bengali. on two retrieval tasks. We hope to incorporate more domains and languages in the future version to provide an increasingly comprehensive evaluation benchmark for community developers. The specifications of all datasets in AIR-Bench 24.05 are available in Table 17, Table 18, Table 19, Table 20, and Table 21. More details are available in Appendix B.1.
3.2 Diversity Analysis
To analyze the query type diversity of AIR-Bench, we utilize GPT-4o777 gpt-4o-2024-08-06 Achiam et al. (2023) as labeler to label the type of the generated queries. Specifically, given a query, we prompt GPT-4o to select the most suitable type for the query from the optional types. The statistics are grouped by tasks and splits in Table 3. Based on the results, we can make the following observations. Firstly, both the QA and Long-Doc tasks have the highest frequency of what queries, followed by claim queries as the second most common, and how queries as the third. Additionally, the QA task exhibits a more balanced distribution of the other query types, whereas the Long-Doc task shows a lower frequency of when queries and where queries. Lastly, a small number of queries are classified as others, reflecting the diverse types of queries present in AIR-Bench to some extent. Further diversity analysis of AIR-Bench is presented in Appendix B.3.
3.3 Positioning of AIR-Bench
We analyze the positioning of AIR-Bench in this section to highlight extra values of AIR-Bench over existing benchmarks. Firstly, as a diverse and continually evolving benchmark, AIR-Bench enables comprehensive evaluation of existing retrievers while addressing the saturation issue that many popular benchmarks (e.g., MTEB Muennighoff et al. (2023) / C-MTEB Xiao et al. (2024)) face due to intensive in-domain fine-tuning. Furthermore, as an automated evaluation toolkit, AIR-Bench supports ad-hoc evaluations for emerging domain-specific retrieval applications. We also provide experimental results in Section 4.2.
#corpus #queries #positives
R-MSMARCO 8,841,823 6,980 7,437
G-MSMARCO 8,872,840 6,319 31,447
      w/o quality control 8,878,865 7,429 7,429
@@ 4 Experiment
RQ2: What additional evaluation functionalities does AIR-Bench offer compared to MTEB/BEIR?
@@ 4.2 Comparison with MTEB/BEIR (RQ2)
Main Results. As presented in Table 6, we can make the following observations based on the comparison results. 1) LLM-based embedding models generally outperform large-size embedding models on both AIR-Bench and MTEB/BEIR, largely due to the superior generalization ability of LLMs. Besides, BM25 performs worse than all embedding models on both AIR-Bench and BEIR. 2) The QA task and the Long-Doc task in AIR-Bench exhibit a level of heterogeneity. The Spearman rank correlation coefficient between the rankings of the nine LLM-based embedding models across the two tasks is only 0.6, with a p-value of 0.0876. Moreover, as a large-size embedding model, e5-large-v2 even outperforms some LLM-based embedding models on the Long-Doc task. 3) By comparing the results on AIR-Bench and MTEB/BEIR, we observe that better performance on MTEB/BEIR may not indicate better performance on AIR-Bench. For example, according to Li et al. (2024), bge-en-icl utilizes more in-domain training data in MTEB/BEIR than bge-en-icl-e5data and achieves more superior performance on MTEB/BEIR. However, compared to bge-en-icl-e5data, bge-en-icl shows performance degradation on AIR-Bench, including both the QA task (54.46 $\rightarrow$ 53.60) and the Long-Doc task (73.43 $\rightarrow$ 72.62). This suggests that increased in-domain training data in MTEB/BEIR may lead to over-fitting, thereby reducing the generalization ability of embedding models.
@@ 5.1 Evaluation Datasets for IR
In recent years, a series of milestone works have been introduced to the community. As the earlier contributions, MS MARCO Bajaj et al. (2016) includes Bing search questions paired with human-labeled relevant passages from Web documents. Natural Questions (NQ) Kwiatkowski et al. (2019) consists of Google search queries with human-labeled relevant Wikipedia pages. Both MS MARCO and NQ are designed for open-domain question answering tasks in English. Recent works like Mr.TyDi Zhang et al. (2021) and MIRACL Zhang et al. (2023) focus on multilingual retrieval in non-English languages. Mr.TyDi covers 11 languages and MIRACL encompasses an extended 18 languages. BEIR Thakur et al. (2021) and MTEB Muennighoff et al. (2023) are introduced to benchmark IR models in a general-domain zero-shot setting, including multiple existing datasets from diverse tasks and domains.
@@ 6 Conclusion
While AIR-Bench aims to be a comprehensive IR benchmark by introducing new features to address the limitations of existing benchmarks, it still has several inherent constraints: 1) Dependence on real-world corpora. The dataset generation process in AIR-Bench begins with corpus preparation. As a result, access to real-world corpora is essential for constructing evaluation datasets. Fortunately, this requirement is typically both feasible and practical in real-world scenarios. 2) Reliance on capabilities of LLM. The quality of the generated testing data in AIR-Bench largely depends on the LLM’s capabilities. However, This limitation can be mitigated by the rapid advancement of LLMs. 3) Potential biases from quality control models. In addition to the LLM, we incorporate several existing IR models during the quality control stage. This reliance may introduce potential biases into the final evaluation datasets. However, as these models continue to improve, the impact of such biases can be progressively reduced.
@@
Acknowledgements
This work is supported by National Science and Technology Major Project (2023ZD0121504), National Natural Science Foundation of China (No. U24A20253, 62276171), Guangdong Basic and Applied Basic Research Foundation (No. 2024A1515011938), Shenzhen Fundamental Research-General Project, China under Grant (No. JCYJ20240813141503005). We appreciate the valuable feedback from Tom Aarsen, Niklas Muennighoff, Jiajun Wang, and Linpeng Tang.
@@ Appendix A Details on Benchmark Construction
In this section, we provide more details on the construction of datasets in AIR-Bench.
@@ A.1 Corpora Preparation
AIR-Bench currently covers two different tasks: QA and Long-Doc. For QA task, we directly use the real-world dataset as the corpus, such as Wikipedia, mC4Raffel et al. (2020), CC-NewsHamborg et al. (2017), etc. We filter out text that is either too short or too long and make a straightforward attempt to remove any information that names or uniquely identifies individuals, as well as any offensive content. For Long-Doc task, we first select one long document for each dataset, such as book, ArXiv paper, legal document, etc., and remove table of contents and references. Then we use the node parser [^13] SimpleNodeParser: https://github.com/run-llama/llama_index tool from LlamaIndexLiu (2022) to split the long document into fixed-size chunks141414chunk_size=200, chunk_overlap=50 as the corpus. All corpora used in AIR-Bench are available in Appendix B.1.
@@ A.3 Quality Control
Embedding Model. Considering that the corpora in AIR-Bench are multilingual, we use bge-m3151515https://huggingface.co/BAAI/bge-m3 as the embedding model to recall the top-1000 relevant documents.
Multiple Re-ranking Models. For the datasets in English and Chinese, we use the following three re-ranking models: bge-reranker-large [^16] https://huggingface.co/BAAI/bge-reranker-large, bce-reranker-base_v1171717https://huggingface.co/maidalun1020/bce-reranker-base_v1, mmarco-mMiniLMv2-L12-H384-v1181818https://huggingface.co/nreimers/mmarco-mMiniLMv2-L12-H384-v1. For the datasets in the other languages, we use the following three re-ranking models: bge-reranker-v2-m3191919https://huggingface.co/BAAI/bge-reranker-v2-m3, mmarco-mMiniLMv2-L12-H384-v1, bge-reranker-v2-gemma [^20] https://huggingface.co/BAAI/bge-reranker-v2-gemma.
@@ 24.05 Oct 17, 2024 9 13 69 Table 17-21
For each dataset, we use the same format as BEIR, i.e. corpus, queries and qrels212121qrels are the relevance labels for queries. The relevance label is 1 for the positive document, and 0 for the negative document., which are all available in the Hugging Face Hub [^22] https://huggingface.co/AIR-Bench of AIR-Bench. To avoid the possible data leakage, we keep the qrels in test splits private. For the qrels in dev splits, we make them public to enable the developers to perform evaluation by themselves.
@@ B.2 Licenses
In Table 16-21, we also list the licenses of the source corpora used for the dataset generation in AIR-Bench. All generated testing data in AIR-Bench is licensed under CC BY-NC-SA-4.0252525https://creativecommons.org/licenses/by-nc-sa/4.0. The testing data in AIR-Bench may only be used for evaluation purposes.
@@ B.3.1 Query Diversity
We also analyze the style diversity of the generated queries in AIR-Bench. We still utilize GPT-4o262626 gpt-4o-2024-08-06: https://platform.openai.com/docs/models/gpt-4o as labeler to label the style of the queries in AIR-Bench. The optional query styles include: formal, informal, professional, casual, complicated, concise, academic, and others. The statistics are grouped by tasks and splits in Table 10.
@@ Appendix C AIR-Bench Software
Appendix D AIR-Bench Data Examples